OpenStudy (anonymous):

multiple choice, Find the slope of the line that passes through the points (1, 10), and (-1, 5) 1)2 1/2 2)-5/2 3)2/5 4)-2/5

OpenStudy (anonymous):

Slope is the "rise over the run." The y co-ordinates in each point can be subtracted to find the rise. The x co-ordinates in each point can be subtracted to find the run. It doesn't matter which co-ordinate you subtract, but you must subtract the same way for both the x and the y co-ordinates. rise = 10 - 5 = 5 run = 1 - (-1) = 2 So, the slope is 5/2 or 2 1/2 as a mixed number. You could also find the same result: rise = 5 - 10 = -5 run = (-1) -1 = -2 So, the slope is -5/-2 or 5/2 or 2 1/2.